### v3

# Sea ice floe segmentation in close-range optical imagery using active contour and foundation models

Abstract
The size of sea ice floes is a key factor influencing ice coverage, albedo, wave propagation through ice-covered waters, and ocean-atmosphere energy exchanges. Floe size can be observed by processing visual-range imagery from ships, aircraft, or satellites. However, autonomously capturing floe boundaries from imagery remains challenging, particularly due to the heterogeneity of sea ice, which impairs boundary definition and reduces image clarity. This study evaluates the accuracy of sea ice floe segmentation using the gradient vector flow (GVF) algorithm, the deep learning-based Segment Anything Model (SAM), and a hybrid approach combining GVF and SAM. These methods are applied to the most extensive dataset of close-range, high-resolution sea ice imagery currently available, collected from cameras aboard an icebreaker during an Antarctic winter expedition. Spanning a wide range of ice conditions and image clarity, the dataset provides a rigorous test of segmentation approaches. Their performance is assessed in terms of floe detection accuracy, floe size distribution, and ice concentration, with results compared against a manually segmented benchmark. The findings indicate that SAM, when used in prompt-driven mode, offers the best balance between accuracy and computational efficiency. Its strong performance in estimating sea ice concentration and detecting floes, while maintaining close agreement with benchmark floe size distributions, makes it suitable for real-time applications and scalable analysis of large imagery datasets.
Keywords sea ice  $\cdot$ ice floe identification  $\cdot$ computer vision  $\cdot$ image segmentation  $\cdot$ active contour model  $\cdot$ foundation model  $\cdot$ segment anything model
1 Introduction
Sea ice modulates the interaction between the atmosphere and ocean at high latitudes (e.g., Landwehr et al., 2021; Bennetts et al., 2024). Its role in the marginal ice zone (MIZ), a dynamic transition region between open ocean and pack ice, is particularly complex and not yet fully understood due to sea ice heterogeneity (e.g., Alberello et al., 2022; Tersigni et al., 2023). The MIZ is characterized by a mix of fragmented, highly mobile sea ice pieces called floes (Alberello et al., 2022; Vichi, 2022; Tersigni et al., 2023). These floes vary widely in type, concentration, and size and are continuously reshaped by ocean currents, waves, and atmospheric and thermodynamic processes (e.g., Herman et al., 2021; Tersigni et al., 2023). The fragmentation of sea ice governs how waves transmit energy through it and, concurrently, how ice further breaks up under their forcing (Dolatshah et al., 2018; Montiel et al., 2016; Meylan et al., 2021; Passerotti et al., 2022; Day et al., 2024). It also influences the ocean surface energy budget by modulating heat fluxes (Andreas et al., 2010; Bourassa et al., 2013; Tersigni et al., 2023), thereby affecting melting processes (e.g., Horvat et al., 2016). Moreover, the floe size distribution influences sea ice rheology, controlling ice deformation and movement under external forces. This, in turn, affects overall ice dynamics, including the formation of ridges, leads (narrow openings in sea ice), and regions of open water (Feltham, 2005; Alberello et al., 2020; Herman et al., 2021; Womack et al., 2022, 2024), adding further complexity to the MIZ (Tersigni et al., 2023).
Observations of floe geometry, coverage, size, and dynamics are key to advancing and validating sea ice models (e.g., Buckley et al., 2024; Roach et al., 2018). Their characterization can be inferred from high-resolution imagery acquired via satellite, airborne, and shipborne sensors, combined with image processing techniques (e.g., Alberello et al., 2022; Alsharay et al., 2022; Herman et al., 2021; Lu et al., 2016; Paget et al., 2001; Parmiggiani et al., 2018; Sandru et al., 2020; Toyota et al., 2006; Zhang and Hughes, 2023; Zhang and Skjetne, 2015; Weissling et al., 2009). Traditional methods, such as threshold segmentation, rely on user-defined values, which can introduce inconsistencies due to variations in sea ice conditions across the image (Heyn et al., 2017; Lu et al., 2008; Alberello et al., 2019). To address this limitation, more advanced techniques, including k-means clustering (Weissling et al., 2009; Lu et al., 2016), the gradient vector flow (GVF) snake algorithm (Zhang and Skjetne, 2015), and watershed segmentation (Dumas-Lefebvre and Dumont, 2023; Parmiggiani et al., 2018), have been applied to improve automation and robustness in sea ice analysis. Despite their advantages, these methods face challenges when applied to complex atmospheric and ice conditions that reduce the clarity of sea ice imagery. Although they perform well under optimal conditions, accuracy and consistency decline markedly in images with reduced clarity. Factors such as coarse resolution, cloud cover, fog, variable lighting, and low contrast between ice and water commonly reduce image clarity. Additionally, snow cover and the presence of interstitial ice, thin ice forming in the spaces between floes, can obscure floe boundaries, creating the illusion of a more uniform surface or larger floes. These challenges complicate the identification of individual floes, particularly when they are densely packed, and make automated analysis of large image datasets particularly problematic.
In recent years, deep learning has significantly advanced image processing, demonstrating high accuracy in object detection, classification, and segmentation across domains such as medical imaging, autonomous vehicles, and environmental monitoring. Convolutional neural networks (CNNs) are commonly used to classify pixels based on patterns learned from training data. In the context of sea ice, CNNs have been applied to differentiate elements such as ice, water, and sky (Zhang et al., 2022; Panchi et al., 2021; Dowden et al., 2021), prioritizing semantic segmentation rather than distinguishing individual ice floe properties. Applications specifically aimed at floe characterization have shown promising results (e.g., Zhang and Hughes, 2023), though over-segmentation remains a challenge, particularly in more heterogeneous sea ice conditions. This is partly due to the reliance of deep learning models on large labeled datasets (Maška et al., 2023). In this respect, the lack of sufficiently diverse training data that accurately captures the complexity and variability of sea ice remains a significant hurdle.
Vision foundation models have emerged as a promising alternative for image segmentation tasks (Dosovitskiy et al., 2020). These large-scale deep learning models are pre-trained on massive datasets and can be adapted to new tasks with minimal additional training, making them particularly suitable for applications where labeled data is scarce (Raghu et al., 2021). One notable example is the Segment Anything Model (SAM), developed by Meta AI Research (Kirillov et al., 2023). Trained on an extensive dataset of one billion masks across 11 million images, SAM can generate segmentation masks without requiring additional labeled data. Its potential for ice observations was demonstrated by Shankar et al. (2023), who applied SAM to segment large icebergs in satellite images, achieving accuracy comparable to specialized CNN models. Nevertheless, SAM was tested on a limited imagery dataset, and its performance in more challenging conditions remains unclear.
Here, we use an image acquisition protocol deployed on the icebreaker S.A. Agulhas II during expeditions to the Antarctic marginal ice zone (MIZ) in the austral winter (Alberello et al., 2019, 2022; Toffoli et al., 2024). The resulting dataset, the most extensive currently available, comprises high-resolution, close-range images of sea ice that offer detailed spatial and temporal coverage, capturing a broad range of ice conditions under varying levels of clarity. Taking advantage of this diversity, we assess the robustness of different segmentation approaches, namely the traditional active contours model (GVF Snake algorithm), the advanced SAM deep learning model, and a novel integration of the two. To support this evaluation, we develop a fully automated processing pipeline that enables their systematic and autonomous application across the varied ice conditions represented in the dataset. We conduct detailed, floe-by-floe visual comparisons against manually segmented benchmarks for two representative scenes with contrasting ice conditions, revealing the segmentation discrepancies, strengths, and limitations of each method. Beyond qualitative assessment, we perform a statistical evaluation using a benchmark of over 2,800 manually segmented floes, the most extensive dataset of its kind for sea ice segmentation analysis, to identify the most effective method for detecting ice floes and estimating floe size and ice concentration under diverse conditions.
2 Image acquisition
Figure 1: Antarctic SCALE expedition, winter 2022: (a) camera mounted on S.A. Agulhas II; (b) ship track from departure on 11 July 2022 to return on 31 July 2022, entering the sea ice region on 19 July and exiting on 24 July; (c–e) high-resolution images of the sea ice surface captured under (c) conditions featuring distinct ice floes with defined edges, (d) diffuse sea ice boundaries characterized by densely packed, irregularly shaped floes with indistinct or visually merged edges, and (e) at nighttime.
Close-range optical sea ice imagery was acquired underway from a sensor installed on the icebreaker S.A. Agulhas II during austral winter expeditions to the Antarctic MIZ in the Eastern Weddell Sea (Fig. 1). These expeditions were part of the Southern oCean seAsonaL Experiment (SCALE). For this study, we focus on data collected during the July–August 2022 expedition (Vichi, 2023). The MIZ was reached at approximately 58S and 1W on 19 July 2022 (Fig. 1b). The vessel then continued southward until it reached consolidated pack ice at latitude 58.8S ($\approx$85 km from the ice edge), marking the inner boundary of the MIZ. Overall, the expedition spent six days in the MIZ before heading back North.
The sensor consisted of a GigE monochrome industrial camera (Fig. 1a) equipped with 2/3” Sony CMOS Pregius sensor and 5 mm F1.8 C-mount lenses (angle of view $\approx$120). It was mounted on the port side of the monkey bridge at 25 m above the water line and tilted by $\approx$70 relative to the ground. The field of view covered a portion of the ocean surface of $\approx$ 200$\times$200 m2. Images were recorded continuously with a resolution of 1920$\times$1080 pixels and at a sampling rate of 1 Hz. The sensor was paired with a synchronized inertial measurement unit (IMU) to record variations in the camera’s position relative to the horizontal plane during navigation.
Around 90,000 images were collected during the expedition, covering both day and night, with nighttime imaging supported by flashlights illuminating the sea ice (see sample images in Fig. 1c–e). Consequently, the database spans a wide range of ice and atmospheric conditions, with approximately 70% capturing complex sea ice scenarios. These include low light conditions (Fig. 1e) and densely packed, irregularly shaped floes with diffuse boundaries, where floe edges are indistinct or merge visually due to snow cover, interstitial ice, or low contrast (Fig. 1d). This diversity offers a robust testbed for evaluating image processing methods.
3 Pre-processing
3.1 Orthorectification
Figure 2: Sample sea ice image (a) from the Antarctic SCALE expedition, winter 2022, after perspective correction, highlighting the region of interest for analysis in red (b). The selected region after image enhancement, including gray-level homogenization through Gaussian bilateral filtering, anisotropic diffusion, and contrast enhancement using CLAHE (c).
The tilt of the sensor causes perspective distortion, making ice floes closer to the camera appear larger than those farther away. This distortion creates inconsistencies in scaling, which vary across each individual image and compromise the accuracy of the conversion between pixel measurements and real-world distances. To address this issue, a correction process known as orthorectification was applied (Luhmann et al., 2023). This process rectifies the sensor orientation by adjusting the distorted pixels as if the image were projected onto a plane perpendicular to the optical axis.
Orthorectification interpolates over the pixels in the original image based on the camera projection matrix, which maps the conversion of a three-dimensional point cloud in the real world to a two-dimensional plane (the image) through intrinsic and extrinsic parameters (Hartley and Zisserman, 2004). The former depends on how the sensor captures images and include, but are not limited to, the focal length, aperture, resolution, and the optical center of the camera’s sensor. These parameters are derived from a calibration process (e.g., Zhang, 2000), which evaluates image distortion by analyzing multiple images of a planar pattern with known geometry (e.g., a chessboard; see Sturm and Maybank, 1999; Bouguet, 2004). Calibration was performed before and after the expedition to verify that intrinsic features remain constant throughout the campaign.
The extrinsic parameters, on the other hand, describe the translations and rotations of the sensor relative to a reference Cartesian coordinate system. For simplicity, the camera was positioned at the origin of the horizontal axis, so translations in the $x$ and $y$ directions were zero, and at a height of 25 m above the water line. The rotations involve three components: tilt around the transverse axis, which was imposed during deployment; roll around the optical axis, which was minimized during installation to ensure alignment with the horizon; and heading around the vertical axis, which aligned with the sensor’s direction. The latter is relevant only when targeting a specific object within the field of view, but becomes negligible when the target is an extended area (such as the ocean surface in this case). In this analysis, translations and rotations are assumed to remain constant over time, effectively neglecting the influence of wave-induced ship motion. This approach aligns with similar applications on mobile platforms, such as aircraft (Zhang and Skjetne, 2015; Parmiggiani et al., 2018) and ships (Heyn et al., 2017; Alberello et al., 2019, 2022; Tersigni et al., 2023). A sensitivity analysis addressing ship motion and validating this assumption is discussed in §3.2.
Orthorectification is performed using the CameraTransform Python package (Gerum et al., 2019), while calibration is carried out with MATLAB’s Computer Vision Toolbox (Bouguet, 2004). The process focused on a rectangular portion of the ocean surface, approximately 95 m $\times$ 165 m, selected based on a visual assessment to include clearly visible floes. The resolution (i.e., pixel-to-metre conversion) was set to 0.05 m to ensure both high object accuracy and the ability to capture small-scale sea ice features (in the order of tens of centimetres; Alberello et al., 2019). A sample image and its orthorectified counterpart are shown in Fig.2a,b.
Floe edges lose sharpness with distance from the sensor (see upper band of Fig. 2b). To avoid ambiguity in the far field and, hence, reduce the risk of detecting unrealistic floes, the workable portion of the orthorectified images is restricted to an area of about 60 m $\times$ 120 m close to the sensors (see area within the red rectangle in Fig. 2b).
3.2 Effect of ship motion
Figure 3: Uncertainty analysis of extrinsic camera parameters on average floe diameter, with the floes used in the analysis highlighted in red. Variations include (a) $\pm$ 1 m elevation, (b) $\pm$ 1 tilt, and (c) $\pm$ 2 roll. Floe diameters are normalized against values derived from orthorectification using ship-measured extrinsics. The shaded blue area indicates a 5% uncertainty range.
The assumption that the sensor remains fixed relative to the ocean surface holds true in calm seas but does not apply in open ocean conditions (cf. Alberello et al., 2022). Movement of the supporting platform can alter the camera’s orientation affecting orthorectification and floe size dimensions. To assess sensitivity to ship motion, an image captured in clear sky conditions was used as a proxy. The orthorectification was evaluated by artificially introducing misalignments of $\pm$1 m in elevation, $\pm$1 in tilt, and $\pm$2 in roll relative to the reference configuration established during calm seas. These values were chosen as representative of extreme ship motions observed during the expedition. A few clearly distinguishable floes were manually segmented after orthorectification by outlining their contours using photo editing software for analysis (see Fig. 3).
Figure 3 shows the average variation in floe size, quantified by the diameter relative to the dimension extrapolated from the image with the sensor in its original position, as a function of misalignment. Uncertainties in floe size due to elevation variations are confined to within 5% (Fig.3a). Tilt variations have a slightly greater impact (Fig.3b), but the effect remains just over 5%. Roll variations have the least impact, with floe size changes under 2% (Fig. 3c). It is important to note that the impact of ship motion is assessed by considering elevation, tilt, and roll independently. However, the ship’s design and control systems are engineered to minimize the chance that these motions peak simultaneously (Fossen, 2011). Consequently, their combined effect is typically lower than the sum of their individual extremes, supporting the assumption that the camera’s motion can be neglected.
3.3 Image enhancement
Sea ice imagery captured in natural settings is influenced by varying light, contrast, shadows, and impurities (see Fig. 1c–e). Consequently, ice floe boundaries in the raw images are often unclear, making automated segmentation challenging (cf. Maška et al., 2023). To address this, image enhancement techniques are applied to sharpen the edges by homogenizing gray-scale levels (e.g., Parmiggiani et al., 2018; Alberello et al., 2019). It is accomplished by combining an edge-preserving Gaussian bilateral filter (Tomasi and Manduchi, 1998) and an anisotropic diffusion filter (Perona and Malik, 1990), which smooths regions with low gradients while preserving ice floe boundaries. Further enhancement is achieved by improving contrast with contrast-limited adaptive histogram equalization (CLAHE) (Zuiderveld, 1994), which accentuates boundaries and reveals fine details. Fig. 2b,c illustrates the transition from the original to the enhanced image. Notably, the adaptive nature of this method allows it to tailor adjustments to each image’s specific characteristics, eliminating the need for predefined thresholds.
For night photographs, enhancement was limited to areas illuminated by the vessel’s spotlight. These regions are identified using haze removal techniques (He et al., 2011), which estimate the illumination map by inverting the low-light image, applying dehazing, and reinverting the result. This highlights the spotlight-illuminated areas for further analysis.
4 Image segmentation
4.1 Active contours model
Active contour models (otherwise known as snakes) adjust an evolving curve to align with floe boundaries (Kass et al., 1988). Poor initialization of the contour can lead to uncertain boundaries, while contours set close to the true edges enable faster and more precise convergence (McInerney and Terzopoulos, 1996). To eliminate the need for manual initialization, an automated contour generator is introduced. This tool adapts to various floe shapes, making the segmentation process fully autonomous.
Fig. 4 illustrates the phases in the image segmentation process. The process begins by separating the foreground (i.e., the objects) from the background, converting the enhanced ice image into a binary format (Fig. 4b). This is done by selecting an appropriate threshold from the gray-scale histogram of the entire image. Pixels below the threshold are labeled as non-feature pixels (background), which includes both the open water fraction and interstitial ice that lacks a specific gradient, and assigned a value of zero, while those above the threshold are labeled as feature pixels (foreground), representing floes, and assigned a value of unity (represented by the white islands in Fig. 4b). In cases with many objects and low gray-scale contrast, the histogram lacks clear modes, leading to ambiguities in selecting the threshold. To resolve this, the Otsu method (Otsu, 1979) is employed for automated thresholding. The Otsu method optimally determines the threshold by maximizing the variance between foreground and background pixel intensities. To enhance accuracy, this binarization is applied to small clusters of $12\times{}12$ m2, each containing about 20 floes. The smaller cluster size results in more distinct modes in the gray-scale histogram, improving the effectiveness of the automated thresholding (Byun and Nam, 2021).
Figure 4: Illustration of the GVF Snake algorithm workflow: (a) enhanced sea ice image; (b) seeds (red dots) representing floe centers on the binarized image; (c) floe perimeters (red lines) identified from isolated, single-seed objects; (d) elliptical initial contours (red lines) of floes containing a single seed after morphological erosion; (e) circular initial contours (red lines) for remaining unused seed; (f) floes identified directly from the binarization process (as shown in Fig. 4c); (g) floes detected by the GVF snake algorithm initialized from elliptical contours; (h) floes detected by GVF initialized from circular contours; (i) final binary segmentation with all detected floes.
A distance transformation (Paglieroni, 1992; Maurer et al., 2003) is applied to the binary image to compute a distance map, where each feature pixel is assigned a value corresponding to its distance from the nearest non-feature pixel. Local maxima in this distance map are pixels with values greater than or equal to those of all their immediate neighbors. These maxima serve as seeds, each marking the center of an individual object (Zhang and Skjetne, 2015). To improve accuracy, regional local maxima are considered by dividing the image into sub-regions, noting this approach was previously used for the binarization. If a sub-region is sharp, objects within it are well separated and typically have only one seed. Consequently, segmentation of these isolated objects follows directly from the binarization process. This initial phase of floe detection is illustrated in Fig. 4c,f, which shows the identified seeds and their corresponding segmented objects, respectively.
However, adjacent objects are merged through the process in most sub-regions, thus preventing segmentation through binarization alone. This results in large, irregular white regions containing multiple seeds (see Fig. 4b). In such cases, segmentation is performed using the gradient vector flow (GVF) Snake algorithm (Xu and Prince, 1998), which extends traditional active contour models by reducing dependence on the initial contour. This allows us to process objects with complex shapes efficiently, even in noisy or low-contrast images. Unlike traditional models, the initial contour shape does not constrain the final segmentation from adapting to irregular boundaries.
The adjacent objects are then separated by smoothing their edges through a morphological erosion process (Adams, 1993). The newly formed objects are initialized using ellipses, which approximate the shape of ice floes (Alberello et al., 2019; Passerotti et al., 2022). The initial contour is then chosen such that the second moment of the covariance matrix (a 2$\times$2 matrix that captures the object’s spatial spread and orientation) aligns with that of the object (Mulchrone and Choudhury, 2004) (see Fig. 4d).
In many cases, objects share significant portions of their edges with neighboring objects and cannot be separated successfully using morphological erosion. Instead, segmentation is performed by applying circles as initial contours to the seeds, as the circular shape is suitable for smaller floes (e.g., Alberello et al., 2019). Using the distance transform, radii are chosen so that contours remain within the objects but are not too far from the presumed edges (Fig.4e). Once initial contours were assigned to all seeds, the GVF snake algorithm was applied to expand the contours and identify the object’s boundaries (Fig.4g,h).
The GVF snake algorithm produces anomalies and inaccuracies, which can result in features that do not exist in the original image (“hallucinations”). To quantify how closely the contours resemble a perfect circle or how elongated the features are, coefficients of circularity and eccentricity are used (e.g., Ayoub, 2003; Bottema, 2000; Alberello et al., 2019). By applying high-valued thresholds, spurious features in the form of nearly perfect circles or line segments are removed (circularity >1 and eccentricity $\geq$0.9). Additionally, objects with open boundaries at the image’s edges are excluded from the segmentation.
The final result is obtained by integrating outputs from each phase of the analysis (see Fig.4i). On average, the computational time is approximately 70 s per floe on a 2.3 GHz Intel Core i9 8 core processor, with no GPU acceleration. Computational time is reported per floe, rather than per image, to avoid bias from variability in the number of floes per image. A summary of computational times is reported in Tables 1 and 2.
From the final binary mask, each ice floe is identified as a connected region of pixels using eight-neighborhood connectivity, meaning that two pixels are considered connected if they share an edge or a corner. Once the floes are identified, geometric properties such as area, perimeter, and diameter are derived. The area is determined by counting the total number of pixels in each region. The perimeter is obtained by tracing along the region boundary to measure its continuous outline (Biswas and Hazra, 2018). The diameter is calculated by equating the floe to a circle of the same area, thus representing the diameter of that hypothetical circle. The ice concentration is computed by summing the areas of all floes and dividing by the total image size (excluding any masked-out, empty regions introduced by orthorectification).
4.2 Segment Anything Model
Snake-based models rely on pixel-level features to segment images, where the features include smoothness, roughness, and contrast, which are specific to each image. In contrast, modern AI-driven approaches, exemplified by vision foundation models (Dosovitskiy et al., 2020), leverage deep learning to identify general patterns from extensive training datasets. These models process entire images without prior knowledge of object shapes, thus eliminating the need for predefined contours or manual feature extraction.
The Segment Anything Model (SAM), developed by Meta AI Research (Kirillov et al., 2023), is a vision foundation model designed for general-purpose image segmentation. Built on the PyTorch deep learning framework (Ketkar and Moolayil, 2021), SAM operates in two distinct modes. The first (SAM-A) is an automatic mask generator that segments all identifiable elements in an image without user input, autonomously determining segmentation regions based on its internal parameters. The second mode (SAM-P) is prompt-based, allowing users to guide the segmentation process by specifying regions of interest using points or bounding boxes. This mode is particularly useful when precision and control over the segmented output are essential.
Tests on two representative images indicate that the newer version, SAM 2 (Ravi et al., 2024), performs worse than the original SAM model for sea ice segmentation tasks, detecting on average 40% fewer floes in automatic mode and 13% fewer in prompt-based mode. A detailed comparison between SAM and SAM 2 is provided in §A. This finding contrasts with Ravi et al. (2024), who reported similar performance between the two versions, but aligns with Sengupta et al. (2024), who found that SAM 2 is less effective on low contrast images. Accordingly, we continue with the original SAM model.
4.2.1 Automatic mask generator (SAM-A)
Figure 5: SAM’s automatic segmentation mode (SAM-A): (a) grid-based fixed points automatically selected by SAM on the enhanced sea ice image; (b) consolidation process applied to the list of masks returned by the model after filtering; (c) final unified binary mask containing all detected floes.
SAM-A autonomously segments an image by selecting a set of sample points from a fixed grid that spans the entire input image (see Fig.5a for a visual representation of the grid layout). For each selected point, the model generates one or more masks, where pixels corresponding to identified objects were labeled as feature pixels (value of one), and pixels representing the background were labeled as non-feature pixels (value of zero). SAM processes the masks to remove redundancy by identifying and eliminating overlaps or duplicates. It then filters the remaining masks to ensure quality, retaining only those with high predicted Intersection-over-Union (IoU) scores, indicating the model’s confidence in mask accuracy, and high stability scores, reflecting robustness to minor perturbations (Kirillov et al., 2023). The final output is a list of filtered masks representing all distinct elements in the image that SAM successfully segments (Fig.5b).
SAM cannot differentiate between the various segmented elements and, hence, segmentation includes both ice floes and open water fractions. The latter is excluded by applying the convexity coefficient (Sklansky, 1970) as a filtering criterion. The complexity coefficient quantifies how closely an object’s boundary resembles that of a convex shape. Ice floes, with regular boundaries that can often be approximated by ellipses or circles (Alberello et al., 2019), typically exhibit convexity values close to 1. In contrast, open water regions, with more irregular boundaries, have lower convexity values. A cut-off value of 0.8 is set, such that masks with a convexity less than or equal to 0.8 are excluded (cf. Wang et al., 2024). The remaining masks (floes) are analyzed to identify those with anomalous shapes or open boundaries at the edges of the image using eccentricity and circularity metrics (see §4.1). Masks exhibiting these irregularities are also eliminated because they are either inconsistent with typical floe geometry or incompletely segmented due to edge effects.
Closer examination of the output reveals that some masks represent the same floes multiple times (see the first two layers in Fig.5b), indicating that the model’s de-duplication process is not fully effective. This is common with adjacent or overlapping floes that have poorly defined boundaries. The redundant masks are then autonomously consolidated into a single unified mask (cf. Shankar et al., 2023), through a pixel-wise union operation, where each pixel is marked as a feature pixel in the final binary mask if labeled as such in any initial mask. This process removes redundancy, ensuring that multiple occurrences of the same feature pixels across different masks are counted only once in the final mask (see Fig.5c).
Geometrical properties of floes and sea ice concentration are extracted from the final unified binary mask. The runtime for segmentation and floe analysis is approximately 0.8 s per floe.
4.2.2 Prompt-based (SAM-P)
Figure 6: SAM’s prompt-based segmentation mode (SAM-P): (a) seed points (red dots) marking floe locations on the enhanced sea ice image; (b) consolidation process applied to the list of masks returned by the model after filtering; (c) final unified binary mask containing all detected floes.
In prompt-based mode, segmentation is guided by user-defined seeds, which are-derived from the distance transform matrix introduced in §4.1 (Fig.4b). These seeds pre-identify the floes, and, for each one, SAM-P generates a corresponding mask (Fig.6a). In regions with poorly defined boundaries, the assignment of seeds to floes can be ambiguous, and the resulting mask does not always isolate the intended floe accurately. In such cases, the mask includes unwanted areas, such as discontinuous water regions or parts of adjacent floes (see the first two layers in Fig.6b). Shape-based filters for circularity, eccentricity, and convexity, with user-defined thresholds (see §4.1 and §4.2.1), are applied to remove clusters of feature pixels that do not match the typical shape of ice floes. A union operation is also used to merge clusters representing parts of adjacent floes with the correct floe. This operation combined all the masks in the output list into a single unified mask, automatically consolidating the segmented portions of floes into their appropriate masks (Fig.6c). From this final binary mask, the properties of each individual ice floe and sea ice concentration are extracted.
The runtime is around 0.3 s per floe. Thus, SAM-P is more than two times faster than SAM-A, and it segments about 15% more floes. The prompt-based mode bypasses the model’s internal point selection process, speeding up segmentation. In contrast, the automatic mode uses a fixed number of points, which often misses some floes, leading to fewer segments.
4.3 A combination of SAM and GVF (SAM-GVF)
Our analysis indicates that the GVF algorithm efficiently delineates floes but encounters difficulties with varied gradients in gray-scale intensity due to its reliance on visual characteristics. SAM, on the other hand, can process a wide range of image variations due to its extended training, but it lacks shape integration, making it difficult to segment adjacent floes with weak edges. We propose combining SAM with GVF to compensate for shortcomings of the former by incorporating shape details, an approach conceptually similar to those suggested previously (Chen et al., 2019; Hoang Ngan Le et al., 2020, e.g.,).
SAM-P is applied to generate an initial binary mask (Fig.7b), from which 90% of the floes are detected, on average (Fig.7b and f). Adjacent floes with blurred or unclear edges are captured by SAM as objects but were not segmented due to the lack of distinct boundaries. The GVF model is applied to analyze these floes. Objects connected by only a few pixels are separated using morphological erosion, resulting in newly delineated edges that serve as initial contours for the GVF model (Fig.7d). For objects sharing significant edge portions, circles are used as initial contours (Fig.7e). These contours are then expanded by the GVF algorithm to trace the true boundaries of each object (Fig. 7g-h). Anomalies, such as spurious features with nearly perfect circles or elongated line segments, are removed using circularity and eccentricity coefficients.
The final SAM-GVF output(Fig.7i) is processed to extract floe properties and sea ice concentration. Segmentation and analysis had a runtime of approximately 4 s per floe.
Figure 7: SAM-GVF hybrid segmentation approach: (a) enhanced sea ice image; (b) unified binary mask from SAM-P showing seeds (red dots) of floes; (c) perimeters (red lines) of floes correctly segmented by SAM-P; (d) initial contours (red lines) of floes containing a single seed after morphological erosion; (e) circular initial contours (red lines) for remaining unused seed; (f) floes identified from SAM-P segmentation (as shown in Fig. 7c); (g) floes detected by GVF initialized from initial contours in Fig. 7d; (h) floes detected by GVF initialized from circular contours; (i) final binary segmentation with all detected floes.
5 Comparison against manual segmentation
5.1 Benchmark dataset
A benchmark is created by manually outlining a subset of twenty images from the original dataset (see §B), which span a wide range of visual conditions and varying levels of sea ice clarity. Segmentation of individual floes is performed using image editing software (Pixelmator, though other tools could also be used). Reference data for the number and size of floes, as well as ice concentration, are subsequently derived. Concurrently, the same sample is segmented using the automated segmentation methods (GVF, SAM-A, SAM-P, and the combined SAM-GVF). Over 2,800 floes were identified across the 20 images, providing a solid foundation for benchmarking (Arbelàez et al., 2011; Martin et al., 2001).
5.2 Visual analysis of segmentation performance on sample images
Two sample images representing contrasting ice morphology, one with distinct ice floes and well-defined edges, and another featuring diffuse boundaries due to closely packed or interconnected floes and interstitial ice, are used to assess the different methods against manual segmentation. This comparison illustrates how varying sea ice conditions influence segmentation outcomes and provides insight into model performances (within their respective constraints). These include: missing floes, where floes identified in the manual segmentation are not detected by the automated methods; misinterpreted floes, which are either partially segmented or merged with adjacent floes; and “hallucinations”, referring to false objects where patterns or textures are mistakenly identified as floes.
Figure 8 and Table 1 present the results for the image with distinct floe edges from the benchmark dataset (original photo shown in Fig. 1c). The chosen image has sharp contrast, uniform lighting, and low noise level, reducing ambiguities. Figure 9 and Table 2 report results for the more challenging case (original photo in Fig. 1d), featuring reduced contrast, subtle lighting variations, and interstitial ice. These factors complicate the detection of floe boundaries, making the automated segmentation more prone to merging distinct floes, overlooking small or faint floes, and introducing false floes where weak edges or texture gradients are mistaken for true boundaries.
5.2.1 Distinct ice floes with defined edges
Figure 8: Visual comparison between manual and automated segmentation methods for an image with distinct ice floes and defined edges from the benchmark dataset. The left column displays the original image overlaid with segmented floe contours (red), showing the manual benchmark (a) and automated methods: GVF (b), SAM-A (d), SAM-P (f), and combined SAM-GVF (h). The right column illustrates segmentation differences relative to the manual benchmark (c, e, g, i), highlighting missing floes (green), misinterpreted floes (yellow), and hallucinations (pink).
The manual segmentation identifies 196 floes with an average diameter of 3.9 m and an ice concentration of 57%. The GVF method misses 35 floes, misinterprets 26, and produces 15 hallucinations (Fig.8c), resulting in approximately 40% of floes not being correctly identified. The missing floes have an average diameter of 2.4 m, representing about 4% of the total surface area. Most misinterpretations are due to partial segmentation of relatively large floes (average diameter of 3.4 m), with few instances of adjacent floes being merged. The hallucinations consist of smaller floes, with an average diameter of 1.6 m. Despite these errors, the GVF method appears to capture the general characteristics of the floe population. However, the relatively high number of missing floes and hallucinations introduces uncertainties in the ice concentration when compared to the manual benchmark.
Table 1: Segmentation results for the distinct edges image using the four methods. For each method, the table reports the number of floes (Fl.), mean diameter ($\overline{\mathrm{D}}$) and ice coverage (IC) associated with the three segmentation errors (missing floes, misinterpreted floes, hallucinations), along with the average computational segmentation time per floe.
Distinct edges image
Method Missing Misinterpreted Hallucinations Average CPU time per floe
Fl. $\overline{\mathrm{D}}$ IC Fl. $\overline{\mathrm{D}}$ IC Fl. $\overline{\mathrm{D}}$ IC
GVF 35 2.39 m 4% 26 3.37 m 7% 15 1.56 m 1% 70 s
SAM-A 15 1.95 m 1% 19 5.29 m 10% 1 2.61 m <1% 0.8 s
SAM-P 8 2.49 m 1% 9 4.25 m 3% 10 1.84 m 1% 0.3 s
SAM-GVF 10 2.63 m 1% 4 3.44 m 1% 7 1.77 m <1% 4 s
SAM-A misses 15 floes, misinterprets 19, and generates one hallucination (approximately 18% of the floes; Fig.8e). The missed floes have an average diameter of 2 m, covering about 1% of the image area, resulting in a more consistent ice concentration estimate than GVF for this specific image. Most misinterpretations result from adjacent floes merging, producing larger floes with an average diameter of 5.3 m and, hence, introducing a bias toward larger floe sizes. The single hallucination covers less than 1% of the image area, making its impact on overall sea ice characteristics negligible.
SAM-P misses only eight floes, and, thus, is the best performing method for capturing floes (Fig.8g). The missed floes have an average diameter of 2.5 m and cover 1% of the image area, resulting in an accurate estimate of ice concentration relative to the benchmark. SAM-P misinterprets nine floes and generates ten hallucinations, bringing the total proportion of incorrectly identified floes to less than 15%. The misinterpreted floes have an average diameter of 4.3 m and are mainly due to merged floes, similar to SAM-A. The number of hallucinations is higher than in SAM-A and is attributed to the model’s susceptibility to complex textures when using prompts to identify uncertain regions (Fan et al., 2023). These hallucinations have an average diameter of 1.8 m and cover 1% of the image area, having minimal impact on the overall floe population and sea ice concentration.
SAM-GVF misses ten floes, misinterprets four, and produces seven hallucinations ($\approx 10\%$ of floes). The missing floes have an average diameter of 2.6 m, covering around 1% of the surface area, similar to SAM-A and SAM-P. Misinterpreted floes have an average diameter of 3.4 m and include both partially segmented and merged floes. Hallucinations have an average diameter of 1.8 m and align with those seen in SAM-P.
Overall, SAM-GVF has the lowest percentage of segmentation errors, with uncertainties affecting $<$3% of the image area. SAM-P achieves similar performance ($\approx$6%), while SAM-A and GVF exhibit the highest uncertainty, affecting 11% and 12% of the surface area, respectively.
5.2.2 Diffuse sea ice boundaries
Figure 9: Visual comparison between manual and automated segmentation methods for an image with diffuse sea ice boundaries from the benchmark dataset. The left column displays the original image overlaid with segmented floe contours (red), showing the manual benchmark (a) and automated methods: GVF (b), SAM-A (d), SAM-P (f), and combined SAM-GVF (h). The right column illustrates segmentation differences relative to the manual benchmark (c, e, g, i), highlighting missing floes (green), misinterpreted floes (yellow), and hallucinations (pink).
As expected, segmentation performance declines under more challenging conditions. For the sample image with diffuse boundaries, the manual segmentation detects 250 floes with an average diameter of 3 m and an ice concentration of 55%. The performance of GVF significantly degrades—it misses 104 floes, which is more than twice the number observed in the distinct edges image. Misinterpretations also increase, with 40 floes affected, while the 17 hallucinations is similar to the distinct edges case (Fig.9g). In total, $\approx$70of floes are not interpreted correctly, where the missed floes have an average diameter of 1.8 m and cover approximately 7% of the image area. The misinterpretations have an average diameter of 3 m, and result mostly from partial segmentation of larger floes, leading to a slight bias in favor of larger floes. The hallucinations have an average diameter of 1.1 m and do not substantially affect the overall sea ice characteristics. However, the increased number of missed and misinterpreted floes reduces the reliability of ice concentration estimates in the diffuse boundaries image.
Table 2: Segmentation results for the diffuse boundaries image using the four methods. For each method, the table reports the number of floes (Fl.), mean diameter ($\overline{\mathrm{D}}$) and ice coverage (IC) associated with the three segmentation errors (missing floes, misinterpreted floes, hallucinations), along with the average computational segmentation time per floe.
Diffuse boundaries image
Method Missing Misinterpreted Hallucinations Average CPU time per floe
Fl. $\overline{\mathrm{D}}$ IC Fl. $\overline{\mathrm{D}}$ IC Fl. $\overline{\mathrm{D}}$ IC
GVF 104 1.77 m 7% 40 2.97 m 8% 17 1.10 m <1% 75 s
SAM-A 81 1.30 m 3% 20 4.49 m 8% 10 1.86 m 1% 0.8 s
SAM-P 53 1.18 m 2% 22 3.25 m 5% 16 1.19 m 1% 0.3 s
SAM-GVF 58 1.24 m 2% 20 2.84 m 3% 12 1.46 m 1% 5 s
SAM-A is more robust for the diffuse boundaries image in comparison to GVF, although it shows a significant decline in performance. Approximately 81 floes are missed, which is over four times more than for the distinct edges image. Misinterpretations remain relatively stable at twenty floes, while hallucinations increase to ten, although this is still fewer than GVF (Fig.8e). Overall, around 45% of floes are not captured correctly. The missed floes have an average diameter of 1.3 m, covering about 3% of the image area. While SAM-A misses more floes in the diffuse boundaries image compared to the distinct edges image, it still retains the capacity to provide an accurate estimate of ice concentration. Misinterpretations have an average floe diameter of 4.5 m, and primarily arise from the merging of adjacent floes, continuing SAM-A’s tendency to bias floe size distribution toward larger floes. Despite the increased hallucinations, their overall impact on sea ice characteristics remains limited.
SAM-P also experiences an increase in segmentation errors, missing 53 floes, which is greater than five times more than in the distinct edges sample. The missed floes are small, with an average diameter of 1.2 m and covering only 2% of the total area. Misinterpretations double to 22 instances, involving smaller floes (average diameter of 3.3 m) than those in the distinct edges image. These misinterpretations primarily result from partial segmentation of floes featuring both snow-covered and bare ice surfaces, which subsequently affects the estimated floe size. Hallucinations increase slightly to 16, likely due to the model’s sensitivity to complex textures when using prompts to refine uncertain regions. Approximately 36% of floes are not interpreted correctly.
SAM-GVF shows moderate performance degradation, missing 58 floes, which is a slight increase compared to the distinct edges image. Misinterpretations remain relatively low, with 20 floes affected, while hallucinations increase slightly to 12, which is fewer than SAM-A and GVF (Fig.9g). Similar to SAM-P, $\approx$36of the floes are not captured correctly. The missed floes have an average diameter of 1.2 m, covering about 2% of the image area. Misinterpretations, with an average floe diameter of 2.8 m, mainly result from partial segmentation of larger floes and those with mixed snow and bare ice surfaces, which slightly distorts the floe size distribution. Hallucinations remain small, with an average diameter of 1.5 m, and do not significantly impact the overall sea ice characteristics.
Of the various methods, SAM-P performs best across both the distinct edges and diffuse boundaries images, exhibiting fewer missed floes and relatively stable rates of misinterpretations and hallucinations. SAM-GVF also demonstrates good performance, while SAM-A and GVF show significant deterioration, particularly with a substantial increase in missed floes. Additionally, computational time per floe increases for GVF (and consequently for SAM-GVF) in the diffuse boundaries image, whereas it remains steady in SAM.
5.3 Floe count, size distribution, and sea ice concentration across the entire sample
Figure 10: Comparison of floe size statistics derived from segmentation methods GVF (green), SAM-A (red), SAM-P (pink), and SAM-GVF (purple) against the manual benchmark (orange). (a) Relative error (RE) in the number of detected floes compared to the manual benchmark (horizontal orange line). b) Violin plots of floe size distributions for each method, with the manual benchmark on the left. Squares inside violins represent medians; darker shaded areas indicate quartiles. (c–f) QQ plots comparing deciles (10% to 90%) of floe size distributions from segmentation methods against the manual benchmark. The solid line indicates perfect alignment. (g) Relative error (RE) in ice concentration compared to the manual benchmark (horizontal orange line).
Data extracted from all 20 images in the sample are compared with the benchmark manual segmentation, as illustrated in Fig. 10. The relative error (RE) in estimating floe numbers is presented in Fig.10a. Among the methods evaluated, SAM-P performs the best, slightly underestimating the number of floes by approximately 2%, which highlights its capability to detect floes across images with varying clarity. In contrast, SAM-A and GVF are the least accurate, underestimating the manual count by approximately 28% and 26%, respectively. This reflects the difficulty in detecting floes in complex scenarios, particularly when dealing with unclear or irregular ice formations (cf. Figs. 8 and 9). For SAM-A, the automated initialization is a key limitation, as it prioritizes larger, more distinct floes while often overlooking smaller or poorly defined ones. The combined SAM-GVF approach mitigates some of the uncertainties in floe detection observed in both SAM-A and GVF, leading to a lower underestimation of the manual count (approximately 12%). This aligns with its performance on the distinct edges image, suggesting that, like SAM-P, it remains relatively consistent across varying image conditions.
The floe size distribution is shown in Fig. 10b as violin plots, which display the overall distribution of data along with summary statistics, such as the median (marked inside the violins) and the first and third quartiles (represented by the darker shaded areas). The benchmark size distribution from manual segmentation exhibits a clear bimodal shape, with peaks around diameters of 1.5 m and 3.5 m. Approximately 50% of the floes fall between 1.8 m and 4.3 m, with a median diameter of 3 m. SAM-P aligns well with the benchmark, with the bimodal shape, median, and quartiles generally well captured. However, the distribution is slightly skewed toward smaller floes, as the first quartile underestimates the benchmark (1.5 m instead of 1.8 m), exaggerating the presence of smaller floes and reflecting the relatively large count of hallucinations (Fig. 8g). Both GVF and SAM-GVF methods also capture the overall shape of the distribution, including the median (3.2 m and 3.1 m, respectively) and quartiles. Minor discrepancies arise from the lower detection of smaller floes, resulting in a less pronounced bimodal shape compared to SAM-P. Conversely, SAM-A produces the least accurate distribution, missing the bimodal nature of the benchmark, resulting in a single peak around 3.3 m. It overestimates both the median (3.8 m) and quartiles, reflecting the limitations of SAM-A in detecting individual floes without seed-based prompts.
The comparison between benchmark and model distributions is further illustrated and analyzed in Figs.10d-g, which present the deciles of the distributions as QQ plots. Despite differences in the overall bimodal shape, the deciles of SAM-P, SAM-GVF, and GVF align reasonably well with the benchmark (Figs.10c,e,f). The lowest deciles (10–30%) are slightly more pronounced, especially for SAM-P, suggesting that these methods detect a large number of small floes, attributed to spurious features or noise. However, the higher deciles align well with the benchmark. The root mean squared error (RMSE) of floe diameter is 0.20 m for SAM-P, 0.17 m for SAM-GVF, and 0.21 m for GVF. Given that the pixel-to-meter conversion in the orthorectification process was set to 0.05 m, the RMSE is approximately 3–4 times the smallest measurable unit in the images. SAM-A (Fig. 10d) shows a more significant underestimation compared to the other models, especially in the lowest five deciles. The RMSE for SAM-A is 0.94 m, reflecting its misalignment with the benchmark’s bimodality.
The relative error associated to sea ice concentration, relative to the benchmark, is shown in Fig.10g. Both SAM-A and SAM-P yielded estimates within 3% of the benchmark, with SAM-A performing slightly better. However, the accuracy of SAM-A is ambiguous, as it results from compensating factors, which are the detection of fewer floes (Fig.10a) and the overestimation of their sizes (Fig. 10b). In contrast, SAM-P’s accuracy stems from more consistent detections of both the number and size of floes. The other two methods underestimate concentration. SAM-GVF yields an error of approximately 10%, while GVF alone performs the worst, with a discrepancy of nearly 30%. The errors in GVF are attributed to missing floes and underestimation of their sizes, particularly in areas with inconsistent gradients across the floes.
6 Conclusion
Three segmentation approaches have been evaluated for analyzing close-scale optical imagery of sea ice. The first approach employs the gradient vector flow (GVF) snake algorithm, which uses edge-based energy minimization for contour detection. The second approach utilizes the Segment Anything Model (SAM), a deep learning framework for image segmentation, applied in both its automatic mode (SAM-A) and a prompt-guided approach (SAM-P). The third approach combines SAM-P with the GVF algorithm (SAM-GVF) to leverage their complementary strengths. To ensure a robust and systematic evaluation, we implemented a fully automated image processing pipeline, incorporating orthorectification and image quality enhancement to standardize images for segmentation analysis. The pipeline was developed to autonomously execute the traditional GVF Snake algorithm, and strategically integrate SAM-P through automatically generated prompts directly targeting individual floes. Performance was quantitatively assessed using a benchmark dataset consisting of 20 manually segmented images representing over 2800 floes under varying lighting, weather, and ice conditions. The evaluation highlighted the relative strengths and weaknesses of each method by assessing their ability to identify and segment individual floes, estimate floe size distribution, and compute sea ice concentration. To our knowledge, this work presents the first evaluation of sea ice segmentation methods using a large-scale, manually annotated benchmark, and introduces an automated framework that enables reproducible, scalable, and robust segmentation analyses across diverse sea ice conditions.
The GVF method was found to provide a reasonable floe size distribution due to its ability to capture the shape of floes, but was less effective at detecting all floes, impacting its accuracy in estimating ice concentration. SAM-A was found to perform well on images with distinct floe edges, but its performance deteriorates in more challenging environments. Throughout the sample, SAM-A detects the fewest floes, and its floe size distribution deviates most from the benchmark. Interestingly, it provides an accurate estimate of sea ice concentration, though this result is ambiguous as it arises from compensating factors: fewer floes are detected, but the identified floes are larger. In contrast, SAM-P is more effective at detecting floes and produces a more accurate size distribution. Its performance is consistent across the sample, resulting in a robust estimate of sea ice concentration. The SAM-GVF combination slightly improves SAM-P’s performance by reducing misinterpretation of complex floes, though its overall performance does not notably differ from SAM-P, returning size distributions and sea ice concentration that closely align with the benchmark.
SAM-P and SAM-GVF emerged as the most effective methods, demonstrating strong performance across multiple metrics, particularly in accurately estimating sea ice concentration and improving floe detection, while maintaining good agreement with the benchmark in terms of floe size distribution. SAM-P is computationally fast, making it well-suited for real-time applications, where immediacy is critical, and for processing large imagery datasets to produce statistical properties of sea ice. In contrast, the combined SAM-GVF approach, along with GVF alone, proves more effective at capturing the specific floe size distribution. However, their longer computational time makes them more suitable for data refinement, where the specific information of the shape of floes is demanded.
Open Research Section
The 20 sample images used in this study, along with the image processing scripts, are available on GitHub (Passerotti, 2025).
Acknowledgments
The expeditions were funded by the South African National Antarctic Programme through the National Research Foundation. AA and AT were partially funded by the ACE Foundation and Ferring Pharmaceuticals and the Australian Antarctic Science Program (project 4434). AT and LGB acknowledge supported by the Australia Research Council (DP200102828, DP240100325, and LP210200927). AA acknowledges the EPSRC grant EP/Y02012X/1. MV was supported by the NRF SANAP contract UID118745. We are indebted to Captain Knowledge Bengu and the crew of the SA Agulhas II for their invaluable contribution to data collection. We acknowledge Dr L. Fascette for technical support.
Appendix A SAM vs. SAM 2
To evaluate how different versions of the model affect segmentation performance, we compare the outputs of SAM and the newer SAM 2 using the two representative images: one featuring distinct ice floes with well-defined edges, and another with diffuse sea ice boundaries. For each version, both automatic and prompt-based modes are tested. Manual segmentation serves as a benchmark to assess how accurately each method captures the characteristics of the floe population. A summary of the comparison is provided in Table 3.
Table 3: Comparison of SAM and SAM 2 segmentation results. For each method and image type, the table reports the number of floes (Fl.), mean diameter ($\overline{\mathrm{D}}$), ice concentration (IC), and average computational time per floe. Manual segmentation results serve as reference benchmarks.
Method Distinct edges image Diffuse boundaries image
Fl. $\overline{\mathrm{D}}$ IC Average CPU time per floe Fl. $\overline{\mathrm{D}}$ IC Average CPU time per floe
Manual 196 3.88 m 57% N/A 250 3.05 m 55% N/A
SAM-A 168 4.20 m 55% 0.8 s 166 3.9 m 51% 0.8 s
SAM-P 192 3.88 m 55% 0.3 s 208 3.44 m 53% 0.3 s
SAM2-A 112 4.47 m 40% 0.7 s 86 4.79 m 37% 0.9 s
SAM2-P 175 3.93 m 50% 0.3 s 172 3.74 m 49% 0.3 s
SAM 2 consistently under performs relative to the original SAM across all evaluated metrics. In automatic mode, it detects 33% fewer floes than SAM in the image with distinct edges, and 48% fewer in the image with diffuse boundaries. In prompt-based mode, the reductions are 9% and 17%, respectively.
SAM 2 tends to overestimate floe size, with mean diameters 15% (distinct edges) and 57% (diffuse boundaries) larger than the manual benchmark in automatic mode, and 1% and 23% larger, respectively, in prompt-based mode. In comparison, SAM in automatic mode overestimates floe size by only 8% (distinct edges) and 28% (diffuse boundaries), while its prompt-based mode exactly matches the benchmark in the distinct edges image and shows a modest 12% increase in the diffuse boundaries image.
Ice concentration from SAM 2 is generally lower than that from SAM, especially in the diffuse boundaries image. The average ice concentration errors relative to the manual benchmark are 18% in automatic mode and 7% in prompt-based mode for SAM 2, compared to just 3% and 2%, respectively, for SAM. Computational time per floe remains similar between SAM and SAM 2.
Appendix B Benchmark dataset photos
Figures 11 and 12 present the twenty photos comprising the benchmark dataset, illustrating a range of ice conditions and lighting environments. Approximately 30% of the dataset represents optimal conditions, featuring distinct ice floes with clearly defined edges, while the remaining 70% captures more complex scenarios characterized by diffuse sea ice boundaries.
Figure 11: Ten original photos from the benchmark dataset (out of twenty total), each annotated with timestamps and approximate GPS coordinates.
Figure 12: The remaining ten original photos from the benchmark dataset, each annotated with timestamps and approximate GPS coordinates.